# Bayesian Nonparametrics for Principal Stratification with Continuous Post-Treatment Variables

### v3

@@ Abstract
Bayesian Causal Inference, Principal Stratification, Shared Atom Mixture Model
@@ 1 Introduction
Principal stratification is a well-established causal inference approach that accounts for post-treatment variables (Frangakis and Rubin, 2002). A post-treatment variable is defined as a variable that is potentially affected by treatment, and it is also associated with the primary outcome. Thus, a post-treatment variable can be seen as an intermediate variable between the treatment and the primary outcome of interest. Post-treatment variables may pose inferential challenges for the estimation of treatment effects, but also provide a useful approach to conceptualizing the mechanisms/pathways through which the treatment affects the primary outcome through the intermediate variable.
Principal stratification with respect to the post-treatment variable involves stratifying units into different groups, known as principal strata. Principal strata consist of units that share the same or similar potential values of the post-treatment variable under all possible treatment levels: by definition, principal strata are not affected by the treatment and thus represent an inherent latent characteristic of the units.
Since the assignment of treatment does not influence the principal strata, comparisons of potential primary outcomes at different treatment levels within a principal stratum represent well-defined causal effects (Rubin, 1974). These specific comparisons are referred to as principal causal effects. Researchers have often referred to the dissociative (and associative) effect as the effect of treatment on the primary outcome for the stratum that includes units whose potential post-treatment variables do not change (or change) under different treatment levels (Frangakis and Rubin, 2002; Mealli and Mattei, 2012).
Most of the literature on principal stratification and principal causal effects has focused on the case of binary treatment and the binary post-treatment variable (see, e.g., Angrist et al., 1996). This implies that researchers deal with at most four principal strata that are the possible combinations of values of these two variables (see the broad literature for binary or discrete post-treatment variables, e.g., Cheng and Small, 2006; Imai, 2008; Ding et al., 2011; Mattei and Mealli, 2011; Mealli and Pacini, 2013; Mealli et al., 2016; Jiang et al., 2016, 2022; Mattei et al., 2024). Considering continuous post-treatment variables leads to a potentially infinite number of principal strata, which, in turn, makes inference and interpretation very challenging. Various solutions have been proposed in the causal inference literature. Proposals include: (i) dichotomize the post-treatment variable (Sjölander et al., 2009; Jiang et al., 2022), or model the continuous post-treatment variable conditional on the binary treatment using (ii) fully parametric or semiparametric modeling (Jin and Rubin, 2008; Conlon et al., 2014; Lu et al., 2023; Schwartz et al., 2011), or (iii) nonparametric modeling (Antonelli et al., 2023).
Although these methods provide potential solutions to the challenges of continuous post-treatment variables, they often require strong assumptions or subjective choices. For example, dichotomization of the post-treatment variable requires the choice of the threshold on the post-treatment variable (e.g., Zigler et al., 2012; Antonelli et al., 2023). Although this choice might be informed by the relevant subject matter literature, it entails a level of arbitrariness/subjectivity and requires extensive checking of the robustness of the results with respect to the chosen thresholds.
Our proposed method introduces a novel and more interpretable approach to characterizing the principal strata, addressing some critical challenges in this framework. To do so, we propose a novel Bayesian approach to principal stratification. Specifically, we introduce a Confounders-Aware SHared-atoms BAyesian mixture model (denoted by the acronym CASBAH, with the intentional alteration of the letter “H") endowed with innovative estimands for principal causal effects. CASBAH exploits the dependent Dirichlet process (MacEachern, 1999; Quintana et al., 2022) to establish a highly flexible distribution for the potential post-treatment variable given the confounders and the treatment variable, similar to the recent proposal of Zorzetto et al. (2024) in the context of heterogeneity in the treatment effects.
Our proposed approach improves the current literature on nonparametric modeling for principal stratification in many respects. First, it allows for the inclusion of the confounders in the weights of the infinite Bayesian mixture model, which provides a way to control for measured confounding and a more precise estimation of the causal effects. Secondly, CASBAH, which leverages shared random kernels’ atoms, allows us to identify coarsened associative/dissociative strata borrowing information across the treatment levels. Hence, our approach eliminates the need for subjective criteria to cluster principal strata. Instead, CASBAH simultaneously (i) identifies the principal strata, according to the nature of the observed data, and (ii) imputes the missing post-treatment variables and primary outcomes. Moreover, our framework naturally incorporates uncertainty quantification for principal strata membership while providing robust estimation of principal causal effects.
@@ 3.1 Confounders-Aware Shared-atoms Bayesian Mixture Model
$\text{P}(Y(0),Y(1),P(0),P(1),T,X)=\int_{\Theta}\prod_{i=1}^{n}\text{P}(Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),T_{i},X_{i},\theta)\text{P}(\theta)d\theta,$
@@
For modeling the post-treatment variable, we exploit a dependent nonparametric mixture, following the dependent Dirichlet process approach (MacEachern, 2000; Barrientos et al., 2012; Quintana et al., 2022). In particular, our model also shares some similarities with the hierarchical Dirichlet process (Teh et al., 2004, 2006; Teh and Jordan, 2009) and the recent common atom model of Denti et al. (2023). Specifically, we assume for each $i=1,\dots,n$:
@@
Prior elicitation is completed by assuming for the parameters $\eta_{l}$ and $\sigma_{l}$ in (7)
@@
3.2 Causal Estimands
To define our causal estimands of interest, we postulate that the units can be marginally grouped into clusters with finite first moments with respect to the treatment level. Consequently, each unit $i$ can belong to: (i) the dissociative stratum if $S_{i}^{(1)}=S_{i}^{(0)}$, which is equivalent to $\mathbb{E}_{S_{i}^{(1)}}\{P_{i}(1)\}=\mathbb{E}_{S_{i}^{(0)}}\{P_{i}(0)\}$; or (ii) the associative positive stratum if $S_{i}^{(1)}\not=S_{i}^{(0)}$ and $\mathbb{E}_{S_{i}^{(1)}}\{P_{i}(1)\}>\mathbb{E}_{S_{i}^{(0)}}\{P_{i}(0)\}$ (throughout the paper, we indicate this condition as $S_{i}^{(1)}\succ S_{i}^{(0)}$); or (iii) the associative negative stratum if $S_{i}^{(1)}\not=S_{i}^{(0)}$ and $\mathbb{E}_{S_{i}^{(1)}}\{P_{i}(1)\}<\mathbb{E}_{S_{i}^{(0)}}\{P_{i}(0)\}$ (throughout, we indicate this condition as $S_{i}^{(1)}\prec S_{i}^{(0)}$).
Definition 1 (Causal Estimands)
Conditionally on the strata allocation, we define the causal estimands for the expected dissociative effect and the expected associative effects as
$\tau_{0}=\mathbb{E}\{Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}=S_{i}^{(0)}\},$
$\tau_{+}=\mathbb{E}\{Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}\succ S_{i}^{(0)}\},$
$\tau_{-}=\mathbb{E}\{Y_{i}(1)-Y_{i}(0)\mid S_{i}^{(1)}\prec S_{i}^{(0)}\}.$
These causal estimands are not directly identifiable from the data due to the conditioning on unobservable quantities (i.e., $S_{i}^{(1)},S_{i}^{(0)}$). However, they can be weakly identified by invoking Assumptions (1) and (2).
@@ 3.2 Causal Estimands
Invoking assumptions (1) and (2), the expected value of each potential outcome can be rewritten as the following
$\mathbb{E}\big\{Y_{i}(t)\mid S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big\}$
$\quad\quad\times\text{P}\big(P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big)dp_{0}p_{1}dx;$ (8)
where the inner expectations $\mathbb{E}\big\{Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big\}$ is estimate with the outcome model $Y_{i}\mid T_{i},X_{i},P_{i}(1),P_{i}(0)$, as well the probability $\mbox{pr}\big(P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,S_{i}^{(1)}=s_{1},S_{i}^{(0)}=s_{0},X_{i}=x\big)$, that is defined by the model for the potential post-treatment variables.
The proof is reported in the Supplementary Material.
3.3 Discovery of Principal Strata
While Bayesian nonparametric mixtures excel at capturing cluster structures, our focus shifts to leveraging this framework for causal inference. To estimate our causal estimands of interest, we need to characterize the joint distribution of $\{S_{i}^{(0)},S_{i}^{(1)}\}$ so that we can determine each unit’s allocation to a principal strata. This joint distribution cannot be directly observed due to the fundamental problem of causal inference. Nevertheless, our proposed Bayesian nonparametric model allows for robust estimation of this joint distribution through its hierarchical construction and shared parameters, effectively addressing this problem. In this section, we illustrate how CASBAH enables the characterization of principal strata in a statistically robust and fully data-adaptive manner.
We start from noting that the latent categorical variables $\{S_{i}^{(0)},S_{i}^{(1)}\}$ for each unit $i\in\{1,\ldots,n\}$ serve a dual purpose. First, they determine the allocation probabilities to specific mixture components. Second, they simultaneously define the marginal distributions of the random unit partition based on potential post-treatment outcomes (Quintana, 2006). This formulation creates a natural bridge between mixture model clustering and principal stratification, allowing us to characterize treatment effect heterogeneity through a principled probabilistic structure.
CASBAH is defined such that the atoms $\{\psi_{l}\}_{l\geq 1}$ are shared between the two potential post-treatment outcomes for the same unit. This formulation yields precise posterior probabilities that quantify, for each unit $i$, whether the latent indicators $\{S_{i}^{(0)},S_{i}^{(1)}\}$ map to identical or distinct mixture components. Specifically, for any unit $i\in\{1,\ldots,n\}$, the probability of membership in the dissociative stratum—where $S_{i}^{(0)}\neq S_{i}^{(1)}$—equals the probability that the atom under control differs from the atom under treatment. This probability is provably non-zero and can be derived a priori through the following theorem.
Theorem 2
Given a unit $i\in\{1,\dots,n\}$, with covariates $x_{i}$ following (2), the probability to belong to the dissociative stratum for each unit $i$ is a function of its covariates $x_{i}$, defined as:
$\text{P}\{S_{i}^{(0)}=S_{i}^{(1)}\}=\frac{\rho_{2}(x_{i})[\{1+\rho_{2}(x_{i})-2\rho_{1}(x_{i})\}^{L}-1]}{\rho_{2}(x_{i})-2\rho_{1}(x_{i})}$ (9)
where $\rho_{1}(x_{i})=\mathbb{E}\big[\Phi\big(\alpha(x_{i})\big)\big]$ and $\rho_{2}(x_{i})=\mathbb{E}\big[\Phi\big(\alpha(x_{i})\big)^{2}\big]$ and $0\leq\text{P}\{S_{i}^{(0)}=S_{i}^{(1)}\}\leq 1$.
The proof is reported in the Supplementary Material. The representation of the probability to belong to the dissociative stratum in (9) is further illustrated in Figure 1.
Such probability varies according to the function $\alpha(x)=\beta_{0}+\beta_{1}x$. Thus, the probability of belonging to the dissociative stratum depends on the observed covariates $x_{i}$ for each unit $i$ and also on the posterior distribution of the parameters $\beta=\{\beta_{0},\beta_{1}\}$. The $\beta$ posterior distributions are a function of all observed data, such that the probability of belonging to the dissociative strata is directly characterized from the data.
@@ 3.4 Posterior Inference
Theorem 1 establishes that principal causal estimands are characterized as functions of both the primary response and potential post-treatment outcomes, each involving specific model parameters. Following the Bayesian paradigm, we estimate these parameters through their posterior distributions conditional on the observed data. This approach provides not only point estimates but also complete uncertainty quantification.

### v4

@@ Abstract
Bayesian causal inference, dependent Dirichlet process, principal causal effects, shared atom mixture model.
@@ 1 Introduction
Principal stratification (Frangakis and Rubin, 2002) provides a useful framework for estimating causal effects and investigating causal mechanisms in the presence of post-treatment variables (also referred to as intermediate variables). This framework has become a cornerstone of modern causal inference for problems involving noncompliance, truncation by death, and surrogate endpoints. Central to principal stratification are the concepts of principal strata and the associated principal causal effects. Units are grouped into principal strata according to the relationship between the potential outcomes for the intermediate variable under different levels of the treatment. The principal strata are independent of treatment assignment and, thus, represent an inherent latent characteristic of the units. Within each stratum, comparisons of potential primary outcomes across treatment levels define well-posed causal estimands, known as principal causal effects. In particular, associative and dissociative effects characterize and quantify treatment effects for units whose intermediate responses would change or not under treatment (Frangakis and Rubin, 2002; Mealli and Mattei, 2012).
Most methodological developments in principal stratification have focused on settings in which both treatment and the post-treatment variable are binary (e.g., Angrist et al., 1996). In this case, the number of principal strata is at most four facilitating both interpretation and modeling (see, among many others, Cheng and Small, 2006; Imai, 2008; Ding et al., 2011; Mattei and Mealli, 2011; Mealli and Pacini, 2013; Mealli et al., 2016; Jiang et al., 2016, 2022; Mattei et al., 2024). When the post-treatment variable is continuous, however, the framework induces a potentially infinite collection of principal strata, indexed by the joint potential intermediate responses under treatment and control. This feature fundamentally alters the problem: the strata are latent objects with no natural discretization, rendering classical stratification strategies impractical and complicating both identification and interpretation.
Existing approaches in principal stratification address this difficulty by imposing additional structure. A common strategy is to dichotomize the continuous post-treatment variable (Sjölander et al., 2009; Jiang et al., 2022), thereby reducing the problem back to a finite number of strata. While appealing for its simplicity, dichotomization requires selecting thresholds that are often arbitrary and may affect the conclusions (Zigler et al., 2012; Antonelli et al., 2023). Alternatively, one may specify fully parametric or semiparametric models for the post-treatment variable conditional on treatment (Jin and Rubin, 2008; Conlon et al., 2014; Lu et al., 2023; Schwartz et al., 2011), or adopt nonparametric modeling strategies (Antonelli et al., 2023). Although these approaches provide valuable tools, they typically rely on restrictive structural assumptions or require externally imposed criteria to coarsen the infinite set of latent strata. As a result, the analyst must choose between arbitrariness and rigidity.
1.1 Contributions
In this work, we propose a Bayesian framework that rethinks principal stratification for continuous post-treatment variables. Rather than fixing a priori strata through discretization or parametric assumptions, we treat principal strata as latent subpopulations with common expected values for the post-treatment variable. To implement this idea, we introduce a Confounders-Aware SHared-atoms BAyesian mixture model (CASBAH), built upon a dependent Dirichlet process prior (MacEachern, 1999; Quintana et al., 2022). CASBAH flexibly models the distribution of the potential post-treatment variable under each treatment level, conditional on observed confounders, while employing shared mixture atoms across treatments. Under this formulation, the existence of dissociative stratum has non-null prior probability. This feature is particularly appealing from a causal perspective: it allows, in a probabilistically coherent manner, the possibility that such strata exist in the population, rather than requiring their presence to be imposed deterministically or ruled out a priori as a byproduct of modeling choices.
Our proposed approach advances the literature in several ways. First, by incorporating confounders directly into the mixture weights (similarly to (Zorzetto et al., 2024)), CASBAH enables flexible adjustment for measured confounding within a fully nonparametric specification of the intermediate response distribution. Second, the shared-atoms formulation allows the model to identify coarsened associative and dissociative strata endogenously, eliminating the need for arbitrary thresholding or externally imposed clustering rules. Third, the framework jointly imputes missing potential intermediate responses and outcomes, and propagates uncertainty in principal strata membership through posterior inference, thereby providing coherent uncertainty quantification for principal causal effects. Together, these features yield a flexible and interpretable approach to principal stratification in settings where the intermediate variable is continuous and traditional methods are inadequate.
@@ 2 Causal Inference Setup
Conditional on principal strata, the estimands of interest are the principal causal effects (PCE), such that
$\mathbb{E}\{Y_{i}(1)-Y_{i}(0)\mid g\{P_{i}(1),P_{i}(0)\}\in\mathcal{S}_{s}\},$ (1)
where $g(\cdot)$ is a functional of the potential post-treatment values and $\mathcal{S}_{s}\subseteq\mathbb{R}$ indicates the subset of values that identify each principal stratum $s$.
@@
2.1 Causal Estimands
Following the general definition of causal estimands in eq. (1), we assume that the functional $g(\cdot)$ is the expected value of the difference of the random variable of potential post-treatment for each unit $i$ given the confounders. The partition of $\mathbb{R}$ defining the principal strata is $\mathbb{R}=\bigcup_{s\in\{0,+,-\}}\mathcal{S}_{s}$, with $\mathcal{S}_{+}=\mathbb{R}_{+}$, $\mathcal{S}_{-}=\mathbb{R}_{-}$, and $\mathcal{S}_{0}=\{0\}$ for the associative positive stratum, associative negative stratum, and dissociative stratum, respectively. This assumption is formalized in the following definition.
Definition 1 (Causal Estimands)
We define the causal estimands in eq. (1) for the expected dissociative effect ($\tau_{0}$) and the expected associative effects positive ($\tau_{+}$) and negative ($\tau_{-}$) as
$\tau_{0}=\mathbb{E}\{Y_{i}(1)-Y_{i}(0)\mid\mathbb{E}_{i}\{P_{i}(1)-P_{i}(0)\}=0\},$
$\tau_{+}=\mathbb{E}\{Y_{i}(1)-Y_{i}(0)\mid\mathbb{E}_{i}\{P_{i}(1)-P_{i}(0)\}>0\},$
$\tau_{-}=\mathbb{E}\{Y_{i}(1)-Y_{i}(0)\mid\mathbb{E}_{i}\{P_{i}(1)-P_{i}(0)\}<0\}.$
These causal estimands are not directly identifiable from the data due to the counterfactual outcome of the post-treatment variable and the primary outcome. However, they can be weakly identified by invoking Assumptions (1) and (2).
For clarity, $\mathbb{E}_{i}\{\cdot\}$ denotes the expectation of a random variable for a specific unit $i$, while $\mathbb{E}\{\cdot\mid\mathcal{A}\}$ denotes the conditional expectation taken over units satisfying the condition $\mathcal{A}$.
Proposition 2.1
Under assumptions (1) - (2), we can rewrite the causal estimands as
$\int_{x}$ $\big[\mathbb{E}\{Y_{i}\mid T_{i}=1,g\{P_{i}(1),P_{i}(0)\}\in\mathcal{S}_{s},X_{i}=x\}$-
$\;\mathbb{E}\{Y_{i}\mid T_{i}=0,g\{P_{i}(1),P_{i}(0)\}\in\mathcal{S}_{s},X_{i}=x\}\big]\text{P}(X_{i}=x\mid g\{P_{i}(1),P_{i}(0)\}\in\mathcal{S}_{s})dx;$
where the expected value of the primary outcome can be rewritten as the following
$\mathbb{E}\big\{Y_{i}\mid T_{i}=t,g\{P_{i}(1),P_{i}(0)\}\in\mathcal{S}_{s},X_{i}=x\big\}$
$\quad\quad\quad\times\text{P}\big(P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,g\{P_{i}(1),P_{i}(0)\}\in\mathcal{S}_{s},X_{i}=x\big)dp_{0}p_{1}dx;$ (2)
where inner expectations $\mathbb{E}\big\{Y_{i}\mid T_{i}=t,X_{i}=x,P_{i}(1)=p_{1},P_{i}(0)=p_{0}\big\}$ are estimated with the outcome model $Y_{i}\mid T_{i},X_{i},P_{i}(1),P_{i}(0)$, as well as probability $\mbox{pr}\big(P_{i}(1)=p_{1},P_{i}(0)=p_{0}\mid T_{i}=t,g\{P_{i}(1),P_{i}(0)\}\in\mathcal{S}_{s},X_{i}=x\big)$, which is defined by the model for the potential post-treatment variables.
@@ 3.1 Confounders-Aware Shared-atoms Bayesian Mixture Model
$\text{P}(Y(0),Y(1),P(0),P(1),T,X)=\int_{\Theta}\prod_{i=1}^{n}\text{P}(Y_{i}(0),Y_{i}(1),P_{i}(0),P_{i}(1),T_{i},X_{i}\mid\theta)\text{P}(\theta)d\theta,$
@@
For modeling the post-treatment variable, we exploit a dependent nonparametric mixture, following the dependent Dirichlet process approach (MacEachern, 2000; Barrientos et al., 2012; Quintana et al., 2022). Specifically, we assume for each $i=1,\dots,n$:
@@
The prior specification is completed by assuming
@@
3.2 Discovery of Principal Strata
In the causal inference literature, the definition of the dissociative stratum for continuous post-treatment variables remains a central methodological challenge. When the post-treatment variable is modeled as continuous with a diffuse distribution, the probability that an unit belongs to the dissociative stratum is zero.
Excluding the dichotomization method for post-treatment variables (Sjölander et al., 2009; Jiang et al., 2022), which the authors of this manuscript find very extreme, alternatives in the literature the literature has proposed defining approximate dissociation via ad hoc thresholding rules (Zigler et al., 2012), whereby individuals are assigned to the dissociative stratum if $g\{P_{i}(1),P_{i}(0)\}$ falls into a user-specified cutoff.
In contrast, leveraging the discrete support induced by the Dirichlet process prior, our approach preserves the definition of the dissociative stratum as $g\{P_{i}(1),P_{i}(0)\}\in\mathcal{S}_{0}$ and introduces a fully data-adaptive procedure that assigns a priori a positive probability to belong to this stratum.
Bayesian nonparametric mixtures are well-known to excel at capturing cluster structures, and we tailor them to solve this specific challenge in causal inference framework. Specifically, to estimate our causal estimands of interest, we need to characterize the joint distribution of $\{P_{i}(0),P_{i}(1)\}$, through their latent cluster allocation variable $\{S_{i}^{(0)},S_{i}^{(1)}\}$, so that we can determine each unit’s allocation to a principal strata. This joint distribution cannot be directly observed due to the fundamental problem of causal inference. Nevertheless, our proposed Bayesian nonparametric model allows for robust estimation of this joint distribution through its construction of shared parameters across the two treatment levels. In this section, we illustrate how CASBAH enables the characterization of principal strata in a statistically robust and fully data-adaptive manner.
We start from noting that the latent categorical variables $\{S_{i}^{(0)},S_{i}^{(1)}\}$ for each unit $i\in\{1,\ldots,n\}$ serve a dual purpose. First, they determine the allocation probabilities to specific mixture components. Second, they simultaneously define the marginal distributions of the random unit partition based on potential post-treatment outcomes (Quintana, 2006). This formulation creates a natural bridge between mixture model clustering and principal stratification, allowing us to characterize the heterogeneity of the treatment effect through a principled probabilistic structure.
CASBAH is defined such that the atoms $\{\psi_{l}\}_{l\geq 1}$ are shared between the two potential post-treatment outcomes for the same unit. This formulation yields precise posterior probabilities that quantify, for each unit $i$, whether the latent indicators $\{S_{i}^{(0)},S_{i}^{(1)}\}$ map to identical or distinct components of the mixture. Specifically, for any unit $i\in\{1,\ldots,n\}$, the probability of membership in the dissociative stratum is equals the probability that the atom under control differs from the atom under treatment, i.e. $S_{i}^{(0)}\neq S_{i}^{(1)}$. This probability is provably non-zero and can be derived a priori through the following theorem.
@@ 3.2 Discovery of Principal Strata
Given the model (4), the probability a priori to belong to the dissociative stratum is defined as:
$\text{P}\{S^{(0)}=S^{(1)}\}=\frac{\rho_{2}(x)[\{1+\rho_{2}(x)-2\rho_{1}(x)\}^{L}-1]}{\rho_{2}(x)-2\rho_{1}(x)}$ (10)
where $x$ are the covariates, $\rho_{1}(x)=\mathbb{E}\big[\Phi\big(\alpha(x)\big)\big]$, $\rho_{2}(x)=\mathbb{E}\big[\Phi\big(\alpha(x)\big)^{2}\big]$ and $0\leq\text{P}\{S^{(0)}=S^{(1)}\}\leq 1$.
The proof is reported in the Supplementary Material. The representation of the probability of belonging to the dissociative stratum in (10) is further illustrated in Figure 1.
This probability varies according to the function $\alpha(x)=\beta_{0}+\beta_{1}x$. Thus, the probability of belonging to the dissociative stratum depends on the observed covariates $x$ for each unit $i$ and also on the distribution of the parameters $\beta=\{\beta_{0},\beta_{1}\}$.
The posterior distribution of strata allocation is by definition a function of the likelihood such that the probability is driven by the observed data.
@@ 3.3 Posterior Inference
Theorem 2.1 establishes that principal causal estimands are characterized as functions of both the primary response and potential post-treatment outcomes, each involving specific model parameters. Following the Bayesian paradigm, we estimate these parameters through their posterior distributions conditional on the observed data. This approach provides not only point estimates but also complete uncertainty quantification.